OpenStudy (jacksonjrb):

Need help with a radical equation. Open this post for the equation. Will give medals. Thanks.

OpenStudy (jacksonjrb):

\[\sqrt{2x+9}-\sqrt{x+1}=\sqrt{x+4}\]

OpenStudy (anonymous):

Square both sides of the equation, combine like terms, solve for x

OpenStudy (jacksonjrb):

\[x+8-2\sqrt{2x+9}\sqrt{x+1}=x+4\] @samarasamara What do I do after this?